30 men can paint an estate in 48 hours. after working for 20 hours, 6 men left. how much longer did the painting take?

Answer by richwmiller(17219)

30*48= 1440 man hours needed
30*20 = 600 man hours used
30*48-30*20= 840 more work to be finished
30-6=24 men available
840/24= 35 hours left
It would have taken 28 hours with all 30 men.